SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (01/15/08)

Attendees: Rainer Bartoldus, Mark Tibbetts, Rahmat Rahmat, Debbie Bard, Kim Hojeong.

Meeting Information

 Date      :  January 15th, 2008
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • Yesterday we had a spike in BLT B tracks phi plot. It seems to be induce by beam parameter changes which happened right before this run. Power-cycling crates solved the problem.
  • L1 rate and L3 rate are higher than Y(4S) era. The L1 rate, L3 rate and deadtime for Jan 10th~15th as a function of run number is here. L1 rate is close to 4k Hz, L3 rate is close to 500 Hz, dead time is close to 2%. As long as dead time is under control (not 20~30% order), we'll keep taking data regardless of high L1 rate. L3 rate until 600 Hz should be OK. Rainer will lower the threshold for SingleGamma line, which will not affect L1 rate but will raise L3 rate. They may go in today.
  • DIRC has problem and wondering if it's TRG problem: email from Joe. We noticed it's pure DCT lines which cause it. From their plot, we see about 150 ns shift for bad one from center. It corresponds to about 1 tick. From FastMonPlots, we see the good run has better tick alignment with smaller error. fast mon plot for Run 78259, the bad one, Same for Run 78204, the good one.

EMT

  • EMC ROMs are swapped. Mark will do a calibration and change the reference file accordingly.

DCT

GLT

Online

L3

  • Rainer will put the lowered threshold SingleGamma lines soon, maybe today.

DQM

  • The cosmics reference plot was from a run without SVT and with DCH HV problem. Rahmat changed the JAS reference for cosmics.
  • Y(3S) (Rec. Lumi.)/(mu pair counting) shows 10~15% drop from Run6. It'll be interesting to see the same plot for Y(4S) in Run7.
  • There are difference between R22 and R24. Compare "PS file from ER farm" in R22 dir and same in R24 dir for same run. p.105 track phi shows difference, number of entries are different too.

Trickle injection monitoring

Trigger Configuration

AOB


This page was created by Kim Hojeong; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: Jan 15th, 2008.